# ViCTr: Vital Consistency Transfer for Pathology Aware Image Synthesis

### v2

@@ 8 Training and Implementation Details
Figure 10: Convergence of models.
Fine-tuning. For fine-tuning, we initialized ViCTr Stage-2 with the pre-trained weights from Stage-1 and configured it for the downstream tasks of CT, MRI, and pathological image generation. Fine-tuning was carried out at a $256\times 256$ resolution, using a batch size of 4 with gradient accumulation over 12 steps. The Adam optimizer was used but with a higher initial learning rate of $1\times 10^{-4}$, and a cosine learning rate scheduler for adaptive adjustment throughout training. Fine-tuning was conducted on a 2-node setup, each equipped with 8 Nvidia A100 GPUs 80GB each. Given Table below shows

### v3

@@ Results
B Qualitative
Figure 10: Synthetically generated MRI images using ViCTr.
C Training and Implementation Details
Pre-training. We pre-trained ViCTr Stage-1 using a rectified flow strategy, with the maximum diffusion steps set to 100. The Atlas-8K dataset was used as the foundational dataset, and training was performed at an image resolution of $\times 256$ 256. We employed a batch size of 888, with gradient accumulation over 8 steps. Optimization was carried out using the Adam optimizer with an initial learning rate of $1\times 10^{-5}$, managed by a cosine annealing scheduler to ensure a smooth decay of the learning rate over time. The pre-training phase was conducted on 4 nodes, each equipped with 8 Nvidia A100 GPUs 80GB each, and completed in approximately 52 hours.
Fine-tuning. For fine-tuning, we initialized ViCTr Stage-2 with the pre-trained weights from Stage-1 and configured it for the downstream tasks of CT, MRI, and pathological image generation. Fine-tuning was carried out at a $\times 256$ 256 resolution, using a batch size of 4 with gradient accumulation over 12 steps. The Adam optimizer was used but with a higher initial learning rate of $1\times 10^{-4}$, and a cosine learning rate scheduler for adaptive adjustment throughout training. Fine-tuning was conducted on a 2-node setup, each equipped with 8 Nvidia A100 GPUs 80GB each. Given Table below shows